OpenStudy (anonymous):

Determine if the following function is continous at the point where x = -5. f(x)=a.{ (x-5)/(x^2-25) x is not equal to -5, b.{-10 x=-5

OpenStudy (anonymous):

The letter a{ is (x-5)/((x+5)(x-5)= (x+5) what happens next? letter b is have no X so it stays the same

OpenStudy (anonymous):

NOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO!!!!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

???

jimthompson5910 (jim_thompson5910):

x^2-25 factors to (x-5)(x+5)

jimthompson5910 (jim_thompson5910):

so (x-5)/(x^2-25) (x-5)/((x-5)(x+5)) 1/(x+5) when x = -5, the denominator is x+5 = -5+5 = 0...but you cannot divide by zero So x = -5 is not allowed in the domain causing a point of discontinuity here.